Tarzan Hospital began opcrations on January 1, 20x1, with Following are the ledger accounts at December 31, 20X1: Expenses E2.4. $10,000 of assets, $6,000 of liabilities, and $4,000 of net assets Liabilities 500 Assets 6,000 1,500 500 1,100 3,200 10,000 1,5001,100 5,000 3,200 4,2004,200 Net Assets Revenues 4,000 5,000 Required: Reconstruct the 20X1 transactions of Tarzan Hospital to the extent possible.
Answer #1

​​​​​​1good sold

2)expense recorded 1100+3200=4300

3)assets and Lia w/o

4)assest purchased

5)amount deposited in bank

​​​​​​austna Cash lc D To Sedes al Ben Good Sold) ooo S000 ) Expense A/c Dr|43c+3200) 4300 Being & Re coded SOo To AssestAl sse s Bong Amount DrBy Sales Sood To

Net profit from profit and loss account =700B S じ9袖17000 | Assestl 670D Banh 42.00 C losing Ner s300o Asset-11700-7 Asses t= )1700net assets =assets - Lia

11700-7000=4700